1270.cm of propane at 18.0°C are heated and allowed to expand to 1530.cm through an isobaric process. Calculate the final temperature in °C.

1
Expert's answer
2022-03-06T15:17:16-0500

According to the Charles's law:


"V_1T_2 = V_2T_1"

where "V_1 = 1270cm^3, V_2 = 1530cm^3" are initial and final volumes respectively, and "T_1 = 18\degree C = 291K, T_2" are initial and final temperatures respectively. Expressing "T_2", obtain:


"T_2 = \dfrac{V_2T_1}{V_1} = \dfrac{1530cm^3\cdot 291K}{1270cm^3} \approx 351K = 78.0\degree C"

Answer. "78.0\degree C".